OpenStudy (anonymous):

evaluate the following expression: if a=9, b=6, and c=12

OpenStudy (anonymous):

there isnt one thats all it says

OpenStudy (darthvader2900):

That doesn't make sense.

HanAkoSolo (jamierox4ev3r):

perhaps this is giving you a polynomial

HanAkoSolo (jamierox4ev3r):

remember, when you plug in an expression into the quadratic formula, you use this breakdown of a polynomial: \(\Large ax^2 +bx +c\)

HanAkoSolo (jamierox4ev3r):

so if you plug in the values of a=9, b=6 and c=12, you can get yourself a polynomial

HanAkoSolo (jamierox4ev3r):

@DARTHVADER2900 lol...you weren't even the one who was explaining. Unless you have something to offer, leave this to me please

OpenStudy (darthvader2900):

I was just asking if he/she understood what you were explaining, so please, no need to get mad. lol.

HanAkoSolo (jamierox4ev3r):

lol it's fine. It's just that it seems unnecessary to do that, if I'm the one guiding the user through the steps. It's just logic

OpenStudy (anonymous):

lol

HanAkoSolo (jamierox4ev3r):

lol, sorry about that x'D Are you following thus far? I could help you find the values of x's, if we're just using the general form of a polynomial

OpenStudy (anonymous):

yeah

HanAkoSolo (jamierox4ev3r):

alright, so if we plug the numbers that are given in this problem to the formula that is given, what would the expression be?

OpenStudy (anonymous):

..

HanAkoSolo (jamierox4ev3r):

oh lol, let me clear myself up

HanAkoSolo (jamierox4ev3r):

if a=9, b=6, and c=12 and if you have the formula \(\large ax^2+bx+c\). then you have \(\Large 9x^2 +6x+12\). Make sense so far? @partyrainbow276

OpenStudy (anonymous):

yeah

HanAkoSolo (jamierox4ev3r):

right then. So we can either factor this right here, or we could use the quadratic formula. I'd say that in this case, factoring would be easiest. DO you know how to factor?

OpenStudy (anonymous):

no

HanAkoSolo (jamierox4ev3r):

hmm.. Then perhaps, it may be best if we use the quadratic formula. I could give it to you, if you're not familiar with it

OpenStudy (anonymous):

oh okay

HanAkoSolo (jamierox4ev3r):

right. So the quadratic formula is \(\Huge x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\)

HanAkoSolo (jamierox4ev3r):

for the values of a,b, and c that you were given in this problem, just plug them in to this formula. Then, you can find your values of x

HanAkoSolo (jamierox4ev3r):

So can you do that, or would you like me to show you how it's done?

OpenStudy (anonymous):

show

OpenStudy (tootzrll):

what grade is this?

OpenStudy (tootzrll):

im in 8th grade. it looks like u can just multiply the numbers then average. ya think?

HanAkoSolo (jamierox4ev3r):

alright. since a=9, b=6, c=12, \(\Huge x=\frac{-6\pm\sqrt{6^2-4(9)(12)}}{2(9)}\)

HanAkoSolo (jamierox4ev3r):

@tootzrll It doesn't appear to be so. as we're trying to evaluate an expression.

HanAkoSolo (jamierox4ev3r):

albeit, we don't know what the expression is, so we're assuming that it's polynomial

HanAkoSolo (jamierox4ev3r):

@partyrainbow276 what math are you taking? I find it odd how ambiguous of a question this is

OpenStudy (anonymous):

algebra

OpenStudy (tootzrll):

huh. seems legit. sorry i couldnt help :/

OpenStudy (anonymous):

its fine

HanAkoSolo (jamierox4ev3r):

if you're in algebra, it's highly likely that you will be needing to know the quadratic formula. Would you like some help in simplifying this down?

OpenStudy (anonymous):

yes

HanAkoSolo (jamierox4ev3r):

So we have this: \(\Huge x=\frac{-6\pm\sqrt{6^2-4(9)(12)}}{2(9)}\) we're going to want to start simplifying the inside of the square root first. \(\Huge x=\frac{-6\pm\sqrt{36-432}}{18}\)

HanAkoSolo (jamierox4ev3r):

you with me so far?

OpenStudy (anonymous):

yep

HanAkoSolo (jamierox4ev3r):

okay. So we keep simplifying this down until we have \(\Huge x=\frac{-6\pm\sqrt{-396}}{18}\) This has to be left like this. Therefore, x is equal to two roots. \(\Large x= \frac{-6+396i}{18}\) and \(\Large x= \frac{-6-396i}{18}\)

HanAkoSolo (jamierox4ev3r):

the reason why we have imaginary numbers is because we have the square root of a negative number. Basically, this means that the roots of this polynomial do not cross the x-axis

HanAkoSolo (jamierox4ev3r):

hope this has helped you! :)

OpenStudy (anonymous):

thank youuu

HanAkoSolo (jamierox4ev3r):

you are most welcome